# Beyond Frame-wise Tracking: A Trajectory-based Paradigm for Efficient Point Cloud Tracking

### v1

@@ IV-A Comparison on nuScenes dataset:
To rigorously evaluate our framework under challenging real-world conditions, we conduct comprehensive comparisons on the challenging nuScenes dataset, characterized by complex scenes and sparser point clouds from 32-beam LiDARs data in diverse scenes. As shown in Tab. I, our TrajTrack consistently outperforms all prior trackers across all categories, demonstrating its exceptional scalability and robustness in sparse and diverse environments. For instance, TrajTrack it surpasses a strong baseline, P2P [16], by in the Car and Pedestrian category by 4.13/3.82% in Success/Precision for the Car category and by an even more significant 4.55/6.81% for the Pedestrian category. This robust performance in a sparse environment highlights the core advantage of our trajectory-based paradigm: by leveraging lightweight, long-term motion continuity, our method maintains a stable track where methods relying solely on instantaneous cues falter.
@@
Method P2B[3] PTT[7] BAT[5] LTTR[31]
FPS 40.0 40.0 57.0 23.0
Method M2Track[9] GLT-T [46] PTTR++ [48] M2Track++[10]
FPS 51.2 30.0 43.0 57.0
Method STTracker [17] SeqTrack3D [18] P2P [16] TrajTrack (Ours)
FPS 22.0 38.0 63.9 56.3
Running Speed: Inference speed is also a vital factor for practical applications. We present a comprehensive speed comparison of TrajTrack with other methods in Tab. II. Following common evaluation protocols [3, 6, 16], speed is measured by calculating the average running time of all frames in the Car category. On a single NVIDIA RTX 3090 GPU, TrajTrack achieves 56.3 FPS. Despite the computational overhead incurred by our IMM module, TrajTrack yields significant performance improvements, maintaining a better trade-off between accuracy and speed.
@@
Robustness to Sparsity: We evaluated our method’s performance under varying levels of point cloud sparsity by analyzing its accuracy based on the number of points in the initial template. As shown in Fig. 4, TrajTrack consistently outperforms other leading methods, especially in extremely sparse scenarios. On the nuScenes dataset, where a majority of tracklets begin with fewer than 15 points, our method’s advantage is most pronounced. This highlights its superior ability to leverage long-term motion continuity when instantaneous appearance cues are minimal. To further validate this, we identified a challenging subset where over half the frames contain fewer than 20 points. On these 3,061 tracklets, our method achieves a Success/Precision of 67.28% / 75.69%, significantly outperforming the baseline’s 65.16% / 72.57%, which confirms its enhanced robustness in low-density conditions.
IV-B Ablation Study
@@
MLP-based IMM 67.77 75.67 51.00 81.64
TrajFormer-based IMM 69.28 76.72 51.01 81.89
We analyze the core Implicit Motion Modeling module of our method in Tab.  IV. Compared to the baseline tracker [16], which relies solely on explicit two-frame motion, incorporating a simple MLP-based IMM already yields substantial performance gains, especially for the Pedestrian category (+4.61% Success). This result validates the core premise of our work: leveraging implicit motion continuity from past trajectories is highly effective. By further employing our more powerful TrajFormer-based IMM, the performance is enhanced to its peak. This systematically demonstrates that while the IMM concept itself is beneficial, the TrajFormer architecture is key to fully capturing the complex temporal dependencies in the trajectory data.
@@ V CONCLUSIONS
VI Limitation and Future Work
Despite its strong performance, TrajTrack has several limitations. First, our current Trajectory-guided proposal refinement is a heuristic; a more deeply integrated, learnable fusion mechanism might yield further improvements. Second, the model currently relies only on geometric information. Scenarios with fewer appearance or color features could lead to identity switches during long-term tracking. Third, model latency still needs to improve for on-board application. One promising direction is to use the quantization technique [53, 54] to compress the model and accelerate the inference speed. Beyond trajectory aggregation, we can include the fusion of multi-modal sensor data [49], such as incorporating dense image information alongside LiDAR point clouds. Besides, while our current focus is on LiDAR’s intrinsic geometric cues, advancements in image understanding and generation [55, 56, 57], like diffusion models for cross-domain image composition, demonstrate the powerful capabilities of modern image processing. Future work can investigate how these advanced image features, enriched by generative priors and semantic controls, might contribute to more robust and context-aware 3D tracking, particularly in challenging, sparse conditions.

### v2

@@ IV-A Comparison on nuScenes dataset:
To rigorously evaluate our framework under challenging real-world conditions, we conduct comprehensive comparisons on the challenging nuScenes dataset, characterized by complex scenes and sparser point clouds from 32-beam LiDARs data in diverse scenes. Following the standard protocol for SOT evaluation, we compare exclusively against state-of-the-art SOT methods. As shown in Tab. I, our TrajTrack consistently outperforms all prior trackers across all categories, demonstrating its exceptional scalability and robustness in sparse and diverse environments. For instance, TrajTrack it surpasses a strong baseline, P2P [37], by 2.87/ 2.97% (Success/Precision) in the Car category and by an even larger margin of 1.89/3.70% in the Pedestrian category. This robust performance in a sparse environment highlights the core advantage of our trajectory-based paradigm: by leveraging lightweight, long-term motion continuity, our method maintains a stable track where methods relying solely on instantaneous cues falter.
@@
Method STTracker [6] LTTR[4] GLT-T [36] SeqTrack3D [27]
FPS 22.0 23.0 30.0 38.0
Method P2B[40] PTT[20] PTTR++ [33] M2Track[66]
FPS 40.0 40.0 43.0 51.2
Method TrajTrack (Ours) BAT[65] M2Track++[67] P2P [37]
FPS 54.7 57.0 57.0 63.9
Running Speed: Inference speed is also a vital factor for practical applications. We present a comprehensive speed comparison of TrajTrack with other methods in Tab. II. Following common evaluation protocols [40, 41, 37], speed is measured by calculating the average running time of all frames in the Car category. On a single NVIDIA RTX 3090 GPU, TrajTrack achieves 54.7 FPS. Despite the computational overhead incurred by our IMM module, TrajTrack yields significant performance improvements, maintaining a better trade-off between accuracy and speed.
@@
Robustness to Sparsity: We evaluated our method’s performance under varying levels of point cloud sparsity by analyzing its accuracy based on the number of points in the initial template. As shown in Fig. 4, TrajTrack consistently outperforms other leading methods, especially in extremely sparse scenarios. On the nuScenes dataset, where a majority of tracklets begin with fewer than 15 points, our method’s advantage is most pronounced. This highlights its superior ability to leverage long-term motion continuity when instantaneous appearance cues are minimal. To further validate this, we identified a challenging subset where over half the frames contain fewer than 20 points. On these 3,061 tracklets, our method achieves a Success/Precision of 67.08% / 75.39%, significantly outperforming the baseline’s 65.16% / 72.57%, which confirms its enhanced robustness in low-density conditions.
@@
MLP-based IMM 67.27 74.14 47.68 77.33
TrajFormer-based IMM 68.02 75.87 48.32 78.78
IV-B Ablation Study
Ablation of IMM: We analyze the core Implicit Motion Modeling module of our method in Tab.  IV. Compared to the baseline tracker [37], which relies solely on explicit two-frame motion, incorporating a simple MLP-based IMM already yields substantial performance gains, especially for the Pedestrian category (+2.25% Success). This result validates the core premise of our work: leveraging implicit motion continuity from past trajectories is highly effective. By further employing our more powerful TrajFormer-based IMM, the performance is enhanced to its peak. This systematically demonstrates that while the IMM concept itself is beneficial, the TrajFormer architecture is key to fully capturing the complex temporal dependencies in the trajectory data.
TABLE V: Ablation Study on Different timesteps Values
timesteps Precision (%) Success (%)
H=1H=1(Baseline) 65.15 72.90
$H=2,T=12$ 68.02 75.87
$H=3,T=6$ 67.17 75.29
$H=3,T=12$ 68.07 75.81
$H=4,T=12$ 67.63 75.67
TABLE VI: Ablation Study on Different ${\lambda}_{IoU}$ Values
$\lambda$ Precision (%) Success (%) Latency (ms)
0.20 66.15 74.26 17.77
0.30 66.81 74.83 17.98
0.40 67.46 75.32 18.06
0.50 68.02 75.87 18.27
Hyperparameters: We conducted a hyperparameter ablation on the historical length $H$, prediction horizon $T$, and refinement threshold ${\lambda}_{IoU}$ (Tab. V-VI). The results show that $H=2$, $T=12$ provide the best overall performance, achieving the highest Success score while maintaining competitive Precision. This configuration offers sufficient temporal context without introducing redundant history. In addition, ${\lambda}_{IoU}=0.5$ yields the highest accuracy with only marginal computational overhead, and is therefore adopted as the default setting.
@@ V CONCLUSIONS
Limitations: Despite outperforming baselines in highly non-linear scenarios (e.g., sharp turns, Fig. 8), TrajTrack’s reliance on motion continuity limits full convergence to the ground truth in these extreme cases. This indicates a need for more adaptive motion modeling to handle abrupt kinematic transitions. Furthermore, replacing the current heuristic proposal refinement with a learnable mechanism could yield further improvements. Besides, we can adopt quantization methods [70, 19, 44, 53, 13, 57, 56] to accelerate 3D Tracker for resource-limited application.